Circular No. 9284

Central Bureau for Astronomical Telegrams

Postal Address: Hoffman Lab 209; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbatiau@eps.harvard.edu  ISSN 0081-0304  cbat@iau.org

               URL http://www.cbat.eps.harvard.edu/index.html

Prepared using the Tamkin Foundation Computer Network

V5854 SAGITTARII = NOVA SAGITTARII No. 3 A nova was discovered on July 12.210 UT with the 1.3-m Warsaw

Telescope located at Las Campanas by P. Mroz et al. (as reported at website

URL http://www.astronomerstelegram.org/?read=9246) in the course

of the OGLE-IV survey (cf. Udalski et al. 2015, Acta Astron. 65, 1), with the discovery I-band magnitude given as 14.8 and the position given as α = 17^{h}49^{m}57^{s}.33, δ = −29^{o}14′37′′.9 (equinox J2000.0). Mroz reports the following additional I-band magnitudes for the variable: July 10.197, [20.1; 10.239, [20.1; 12.277, 14.45; 12.320, 14.42; 13.172, 13.91; 14.195, 13.36; 14.269, 13.31.

E. Kazarovets and N. Samus note that the permanent GCVS designation V5854 Sgr has been assigned to this variable. V5853 Sgr (cf. IAUC 9283) is thus Nova Sgr 2016 No. 2.

V407 LUPI = NOVA LUPI 2016 = PNV J15290182–4449409 A new nova (mag 9.1) found on Sept. 24.010 UT via robotic 14-cm tele-

scopes in the course of the V -band All-Sky Automated Survey for Supernova (ASAS-SN) is located at α = 15^{h}29^{m}01^{s}.82, δ = −44^{o}49′40′′.9 (equinox 2000.0). The variable was designated PNV J15290182–4449409 when it was posted at the Central Bureau’s TOCP webpage. PNV J15290182–4449409 brightened to visual magnitude ∼ 6 on Sept. 25–26. Additional details appear on CBET 4322.

Kazarovets and Samus note that the permanent GCVS designation V407 Lup has been assigned to this variable.

V5855 SAGITTARII = TCP J18102829–2729590 As first announced on CBET 4332, S. Nakano, Sumoto, Japan, re-

ported the discovery by Koichi Itagaki of a nova-like object (mag 10.7) on an unfiltered CCD frame (limiting mag 13.0) taken on Oct. 20.383 UT using a 180-mm-f.l. f/4.0 camera lens; a confirming image taken immediately thereafter yielded position α = 18^{h}10^{m}28^{s}.29, δ = −27^{o}29′59′′.0 (equinox 2000.0). The variable was designated TCP J18102829–2729590 when it was posted at the Central Bureau’s TOCP webpage. Nothing was visible at this position on an image taken by Itagaki on Oct. 4.45 (limiting mag 13.0). Visual magnitudes by A. Pearce, Nedlands, W. Australia: Oct. 20.485, 11.0; 21.481, 9.9; 22.479, 8.6; 23.485, 8.0. Spectroscopy by K. Ayani on Oct. 24.38 (cf. CBET 4333) shows this to be nova at early phase.

Kazarovets and Samus note that the permanent GCVS designation V5855 Sgr has been assigned to this variable; this is also N Sgr 2016 No. 4.

2016 December 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2016 CBAT  Daniel W. E. Green